OpenStudy (annipuppi):

|2x+3|=3x+7 show work

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ |2x+3|=3x+7 \rm { }\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 2x+3=3x+7 ~~~~~or~~~~~ 2x+3=-(3x+7)\rm { }\LARGE\color{white}{ \rm │ }}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 2x+3=3x+7 \rm { }\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ -4=x \rm { }\LARGE\color{white}{ \rm │ }}\) OR \(\normalsize\color{blue}{ 2x+3=-(3x+7 )\rm { }\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 2x+3=-3x-7 \rm { }\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ 5x=-10 \rm { }\LARGE\color{white}{ \rm │ }}\) \(\normalsize\color{blue}{ x=-2 \rm { }\LARGE\color{white}{ \rm │ }}\)